# Learning to Animate Images from A Few Videos to Portray Delicate Human Actions

### v1

@@ 3.3 Detail Enhancement Decoder
In LDM, pixel-level details can be distorted when videos are decoded from the latent space, as even slight perturbations within the latent space can lead to noticeable visual artifacts, compromising the intricate details and transition smoothness. To mitigate this issue, we devise the Detail Enhancement Decoder that extends the image decoder $\mathcal{D}$ with additional layers to propagate multi-scale details from the reference image to the generated frames, handling both small and large displacements.
We define the levels of both the encoder and decoder as $l\in\{0,1,\cdots,L\}$, with $l=0$ representing the pixel space and $l=L$ representing the latent space. At level $l$, we extract the decoder features of the $i$-th decoding frame, denoted as $\bm{h}_{l}^{i}$, and the encoder features of the reference image, denoted as $\bm{g}_{l}^{1}$. $\bm{g}_{l}^{1}$ is then propagated to enhance the details in $\bm{h}_{l}^{i}$ through two branches, as shown in Figure 2 (c). The warping branch retrieves details from nearby areas in $\bm{g}_{l}^{1}$ for each spatial position in $\bm{h}_{l}^{i}$. It learns the displacements between the two features and warps $\bm{g}_{l}^{1}$ into the output $\hat{\bm{g}}_{l}^{1}$ based on the learned displacements. The patch attention branch retrieves details from the global scope of $\bm{g}_{l}^{1}$, complementing the local retrieval of the warping branch. It employs an attention layer with $\bm{h}_{l}^{i}$ as the query and $\bm{g}_{l}^{1}$ as the key and value to produce the output $\check{\bm{g}}_{l}^{1}$. The output features from the two branches are fused using learnable weights $\bm{w}^{i}_{l}$: $\tilde{\bm{h}}_{l}^{i}=\bm{h}_{l}^{i}+\bm{w}^{i}_{l}\odot(\hat{\bm{g}}_{l}^{1}+\check{\bm{g}}_{l}^{1})$, where $\odot$ represents element-wise multiplication. The fused features $\tilde{\bm{h}}_{l}^{i}$ is then passed to the next level. Through detail propagation at each level for each decoding frame, the details in the generated videos are enhanced. The Detail Enhancement Decoder is trained to retrieve proper details through reconstructing distorted videos to their ground-truth versions. Details can be found in Appendix Sec. A2.3.
Figure 3: Qualitative comparison of different methods.
@@ 4 Experiments
We conduct experiments on 12 actions selected from HAA500 [7]. The actions include single-person actions (push-up, arm wave, shoot dance, running in place, and sprint run), human-object interactions (soccer shoot, drinking from a cup, balance beam jump, canoeing sprint, chopping wood, and ice bucket challenge), and human-human interactions (hugging human). More details about data and implementation are in Appendix Sec. A3.1 and Sec. A3.2.
@@ 4.1 Main Results
Metrics. Following [69, 70, 23], we use three CLIP-based metrics: Text Alignment, Image Alignment, and Temporal Consistency, in which higher scores indicate better performance. Following [72], we utilize Fréchet distance to compare generated videos and real ones. To mitigate content bias in the commonly used FVD [62], we adopt CD-FVD [12], where a lower distance indicates better performance. To assess the similarity between generated and ground-truth videos in HAA500, we calculate the cosine similarity for each pair of generated and ground-truth videos. We utilize RGB and optical flow to calculate two metrics: Cosine RGB and Cosine Flow. In these metrics, higher similarities reflect better performance. For all metrics, we report the average results across all test videos. More details are described in Appendix Sec. A3.3.
@@
Baselines. We compare FLASH with several image animation baselines, including the zero-shot training-free image animation model TI2V-Zero [48]; large-scale trained models like SparseCtrl [20], PIA [80] and DynamiCrafter [72]; and motion customization models like DreamVideo [68] and LAMP [70]. More details are described in Appendix Sec. A3.4.
@@
Qualitative Results. We compare the qualitative performance of different methods in Figure 3, excluding TI2V-Zero and SparseCtrl due to their lower quality and space limitations. Animated videos of all methods are available on the webpage. Although PIA, and DynamiCrafter are trained on large-scale video datasets, they still generate unrealistic and disjointed motion that diverges considerably from the correct actions. These results reveal the limitations of large-scale pretrained video generative models in animating uncommon human actions. DreamVideo and LAMP finetune video generative models on a small set of videos containing the target actions. While DreamVideo produces realistic actions, it significantly deviates from the reference images. The results indicate that it struggles to adapt motion to different reference images flexibly, because it requires training on each reference image individually. LAMP demonstrates smooth transition from the reference image, but its rendering of the shoot dance displays discontinuities (e.g., disconnected or missing limbs) and it fails to generate the chopping wood action, demonstrating its limitations. In contrast, FLASH not only maintains smooth transition from the reference image but also realistically animates the intended actions that resemble real videos, showcasing its effectiveness.
Table 1: Quantitative comparison of different methods.
Method Text Alignment ($\uparrow$) Image Alignment ($\uparrow$) Temporal Consistency ($\uparrow$)   CD-FVD ($\downarrow$)    Cosine RGB ($\uparrow$)    Cosine Flow ($\uparrow$)
TI2V-Zero 23.30 66.75 87.60 1584.30 0.6859 0.5056
SparseCtrl 21.90 60.77 88.54 1627.87 0.6704 0.5663
PIA 23.13 63.58 93.85 1547.61 0.6958 0.6055
DynamiCrafter 22.60 81.71 95.23 1438.01 0.7980 0.6390
DreamVideo 23.77 64.47 93.47 873.76 0.6672 0.6318
LAMP 22.82 77.93 93.92 1260.46 0.8284 0.6989
FLASH 23.02 79.04 95.64 786.39 0.8626 0.7786
Figure 4: Animated actions generated by FLASH using reference images from the Internet and generated by image generative models.
Table 2: Quantitative ablation studies on different components of FLASH.
Variant Strong Augmentation Motion Features Alignment Inter-frame Correspondence Alignment Detail Enhancement Decoder Text Alignment ($\uparrow$) Image Alignment ($\uparrow$) Temporal Consistency ($\uparrow$)   CD-FVD ($\downarrow$)   Cosine RGB ($\uparrow$)   Cosine Flow ($\uparrow$)
#1 22.53 77.10 95.43 1023.30 0.8380 0.6806
#2 22.48 76.72 94.91 932.92 0.8398 0.7061
#3 22.64 76.48 95.06 920.39 0.8444 0.7140
#4 22.70 76.31 94.84 938.21 0.8432 0.7172
#5 22.52 76.35 95.01 906.31 0.8446 0.7224
#6 22.77 76.22 95.31 908.39 0.8451 0.7233
Figure 5: Qualitative ablation study on different components of FLASH. #1: baseline model, #2: model trained with only strongly augmented videos, #3: with motion feature alignment, #4: with inter-frame correspondence alignment, #5: with both alignments, #6: full model with Motion Alignment Module and Detail Enhancement Decoder. See Table 2 for the correspondence of each variant.
Quantitative Results. We compare FLASH with baselines across six metrics in Table 1. The results show that FLASH achieves the best overall performance, except in Text Alignment and Image Alignment. This suggests that FLASH generates actions with greatest temporal consistency and similarity to real action videos. In terms of Text Alignment, TI2V-Zero and DreamVideo outperform FLASH, but both exhibit significantly lower scores on Image Alignment. This implies that while they can generate correct actions, they struggle to animate reference images to portray specified actions, as shown in Figure 3. In terms of Image Alignment, DynamiCrafter surpasses FLASH, but it performs considerably worse on CD-FVD, Cosine RGB, and Cosine Flow. This indicates that although DynamiCrafter maintains consistency with the reference images, it fails to generate realistic actions, as also observed in Figure 3.
User Study. Given the potential limitations of the CLIP, I3D, and RAFT models, we conducted a user study to further evaluate the quality of the generated videos. This study was conducted on Amazon Mechanical Turk (AMT), where workers were instructed to select the best generated video from a set of candidates. For each action, we randomly select 4 different reference images for evaluation. Control questions were included to identify random clicking, and only answers from workers who correctly answered the control questions were considered valid. More details are described in Appendix Sec. A4.1. Out of 366 valid responses, FLASH was preferred in 67% of the response, significantly outperforming the next best models, DynamiCrafter (14%) and LAMP (12%). These results indicate that FLASH produces videos of the highest quality.
@@
Generalization to Internet and Generated Images. To assess the generalization capability of FLASH beyond the HAA500 dataset, we tested it on images sourced from the Internet and those generated by Stable Diffusion 3 [10]. As shown in Figure 4, FLASH successfully animated a variety of scenes, including a person doing a pushup in an office and running on snow. It also adapted to unrealistic scenarios, such as an astronaut running in place within a virtual space and a cartoon character shooting a soccer ball. Additionally, FLASH can animate generated images, such as a humanoid alien pouring water over his head, two humanoid aliens hugging. More animated videos are available on the webpage. These results highlight FLASH’s strong generalization ability across a broad spectrum of reference images.
@@ 4.2 Ablation Studies
We conducted ablation studies on four actions: sprint run, soccer shoot, canoeing sprint, and hugging human. The quantitative and qualitative results are presented in Table 2 and Figure 5, respectively. Variant #1 serves as the baseline, excluding both the Motion Alignment Module and the Detail Enhancement Decoder. Variant #2 uses only strongly augmented videos without any alignment technique. Variants #3, #4, and #5 progressively incorporate motion feature alignment, inter-frame correspondence relation alignment, and both, respectively. Lastly, Variant #6 builds upon Variant #5 by incorporating the Detail Enhancement Decoder.
@@
Comparing the quantitative results of Variant #2 with Variants #3, #4, and #5, we find that Variants #3, #4, and #5 improve CD-FVD, Cosine RGB, and Cosine Flow. Both Variants #3 and #4 enhance the Cosine RGB, and Cosine Flow. When combined, Variant #5 yields further enhancements in cosine similarity and a 25-point improvements in CD-FVD, without a noticeable drop in CLIP scores. Qualitative results also indicates improved fidelity in Variants #3, #4, and #5. For instance, motion in Variant #2 appears unrealistic in both actions. In the soccer shooting action, the person’s foot didn’t touch the soccer ball, and the leg appears disconnected in some frames. In the canoe paddling action, the hand positions on the paddle are inconsistent across frames. However, these issues are largely mitigated in Variants #3, #4, and #5. These results demonstrate the effectiveness of the Motion Alignment Module in learning accurate motion. By providing explicit guidance for learning appearance-general motion, the module directs the model toward generalizable motion, thereby improving the quality of the generated videos.
Comparing the quantitative results of Variant #5 and Variant #6, we observe that Variant #6 noticeably improves Text Alignment and Temporal Consistency without substantially affecting CD-FVD, Cosine RGB, or Cosine Flow. Qualitatively, Variant #6 enhances some details (e.g., the soccer ball in certain frames in the soccer shooting action) and reduces noise in generated frames. These results suggest that the Detail Enhancement Decoder could compensate for action related details in generated frames, leading to better temporal consistency and alignment with the action descriptions. Since the decoder operates on a frame-by-frame manner without considering inter-frame relations, it has minimal impact on motion patterns, resulting in slight changes on CD-FVD, Cosine RGB, and Cosine Flow.
In Appendix Sec. A4.2, we show that the Motion Alignment Module improves performance across different few-shot settings (i.e., 8 or 4 videos per action class) and benefits from joint training across multiple action classes. Additionally, we provide ablation studies analyzing the effects of hyper-parameters in the Motion Alignment Module and the branches of the Detail Enhancement Module. Limitations are discussed in Appendix Sec. A5.
[…]

### v2

@@ 3.3 Detail Enhancement Decoder
The previous section introduces the Motion Alignment Module, which improves the generalization of the learned motion features in the latent space. However, a strong latent representation per se does not guarantee the quality of the generated video; we still need a powerful decoder that can (1) reproduce natural motion in the decoded video, even though the latent representation may contain distortion or noise, and (2) ensure the later frames in the video flow from the first reference frame smoothly.
To this end, we propose the Detail Enhancement Decoder. We introduce architectural components that directly retrieve multi-scale details from the reference frame and propagate them to later frames. Further, we apply strong data augmentation to the input, so that the decoder must learn to recover from distorted latent representations.
Multi-scale Detail Propagation. We number the network layers in both the encoder $\mathcal{E}$ and decoder $\mathcal{D}$ using $l\in\{0,1,\cdots,L\}$. The features in the encoder are denoted as $\bm{g}_{l}$ and those in the decoder are denoted as $\bm{h}_{l}$. A special case is $\bm{g}_{0}$ and $\bm{h}_{0}$, which respectively denote the original and reconstructed video in pixel space. $\bm{g}_{L}$ is the output from the encoder and $\bm{h}_{L}$ is the input to the decoder; both are in the latent space. We further use superscripts to denote the frame number. $\bm{g}_{l}^{1}$ denotes the encoded feature from the first, user-supplied reference frame, taken from encoder layer $l$. A main goal of the Detail Enhancement Decoder is to propagate this feature to decoder features of later frames.
In addition to the layer-by-layer structure of a typical decoder, we propose two new architectural components, namely the Warping Branch and the Patch Attention Branch. The aim of the Warping Branch is to retrieve relevant appearance information from the reference frame $\bm{g}_{l}^{1}$ for each spatial location $(p,q)$ in the $i$-th frame $\bm{h}_{l}^{i}$. However, due to motion, it is not immediately clear which spatial location in $\bm{g}_{l}^{1}$ is relevant to position $(p,q)$ of the $i$-th frame. We simply apply a neural network $\mathcal{N}$ to predict the displacement $(\Delta p,\Delta q)$. The network takes the two frames $\bm{h}_{l}^{i}$, $\bm{g}_{l}^{1}$ as input, and outputs two scalars for each spatial location,
$(\Delta p,\Delta q)=\mathcal{N}(\bm{h}_{l}^{i},\bm{g}_{l}^{1})[p,q].$ (5)
Hence, the relevant position from the reference frame is $(p+\Delta p,q+\Delta q)$. As $\Delta p$ and $\Delta q$ may not be integers, we perform bilinear interpolation to find the exact feature at that location. We store the retrieved features of all spatial locations in a new tensor $\bm{s}^{i}_{l}$.
To complement the local retrieval of the Warping Branch, we introduce the Patch Attention Branch, which retrieves details from the entire reference feature map $\bm{g}_{l}^{1}$. We divide both $\bm{h}_{l}^{i}$ and $\bm{g}_{l}^{1}$ into patches and apply a standard cross-attention layer $\mathcal{A}$, using $\bm{h}_{l}^{i}$ as the query and $\bm{g}_{l}^{1}$ as both the key and the value. The output features are denoted as $\bm{t}_{l}^{i}$.
We fuse the two output features using weights $\bm{w}^{i}_{l}$ produced by a network $\mathcal{M}$:
$\bm{w}^{i}_{l}=\mathcal{M}(\bm{h}_{l}^{i},\bm{g}_{l}^{1}),$ (6)
$\hat{\bm{h}}_{l}^{i}=\bm{h}_{l}^{i}+\bm{w}^{i}_{l}\odot(\bm{s}_{l}^{i}+\bm{t}_{l}^{i}),$ (7)
where $\odot$ represents element-wise multiplication. We pass the fused features $\hat{\bm{h}}_{l}^{i}$ to the next layer in the decoder, which is layer $l-1$.
Distorted Latent Representation as Input. Given an original video $\bm{X}^{\text{ori}}$, we first distort it into $\bm{X}^{\text{dis}}$ by applying Gaussian blur with a random kernel size, random color adjustments in randomly selected regions, and random elastic transformations. We then encode $\bm{X}^{\text{dis}}$ into a latent video $\bm{Z}^{\text{dis}}_{0}$. This process introduces potential distortions in latent videos. Note that the degree of distortion is randomly sampled, so there are some inputs to the decoder that receive minimal or zero distortion.
Reconstruction Loss. The entire decoder, including the Warping Branch and the Patch Attention Branch, is end-to-end trainable. Therefore, we simply train the decoder to reconstruct $\bm{Z}^{\text{dis}}_{0}$ back to $\bm{X}^{\text{ori}}$. We denote the decoded video from $\bm{Z}^{\text{dis}}_{0}$ as $\hat{\bm{X}}^{\text{ori}}=\mathcal{D}(\bm{Z}^{\text{dis}}_{0})$, and use a reconstruction loss to train the newly added layers (i.e., $\mathcal{N}$, $\mathcal{A}$ and $\mathcal{M}$):
$\mathcal{L}_{R}=\|\hat{\bm{X}}^{\text{ori}}-\bm{X}^{\text{ori}}\|^{2}_{2}.$ (8)
More detail are provided in Appendix A2.3.
@@ 4 Experiments
We conduct experiments on 16 actions selected from the HAA500 dataset [9], including single-person actions (push-up, arm wave, shoot dance, running in place, sprint run, and backflip), human-object interactions (soccer shoot, drinking from a cup, balance beam jump, balance beam spin, canoeing sprint, chopping wood, ice bucket challenge, and basketball hook shot), and human-human interactions (hugging human, face slapping). Most of these classes are challenging for general video generative models to animate. We train a separate model for each action. More details about data and implementation are in Appendices A3.1 and A3.2.
Figure 3: The percentage of users that choose products of each video generator as the best videos in the user study on Amazon Mechanical Turk. The proposed method, FLASH, received the vast majority of votes.
Figure 4: Qualitative comparison of different methods. We strongly recommend watching the animated videos in the Webpage, as motion artifacts are hard to notice from static images.
@@ 4.1 Main Results
We compare FLASH with several baselines, including TI2V-Zero [61], SparseCtrl [25], PIA [106], DynamiCrafter [94], DreamVideo [90], MotionDirector [109] and LAMP [92], whose details are described in Appendix A3.4.
User Study. Existing automatic evaluation metrics fall significantly short of detecting all types of artifacts. Human evaluation remains the gold standard in evaluating the quality of generated videos. Thus, we perform a large-scale user study on Amazon Mechanical Turk to compared the videos generated by FLASH and baselines.
In the user study, workers were tasked to select the video of the highest quality from a set of candidates. For each action, we randomly selected four different reference images and their corresponding generated videos for this user study. To identify random clicking, each question was paired with a control question that has obvious correct answer. The control question includes a real video of a randomly selected action alongside clearly incorrect ones, such as a static video. The main question and the control question were randomly shuffled within each question pair, and each pair was evaluated by 10 different workers. Responses from workers who failed the control questions were marked as invalid. More details are provided in Appendix A4.1. Among 488 valid responses, FLASH was preferred in 65.78% of cases, as shown in Figure 3, which significantly outperforms other methods and highlights its superiority.
@@
Qualitative Results. In Figure 4, we compare FLASH with PIA, DynamiCrafter, DreamVideo, MotionDirector and LAMP, while providing results for other methods on the Webpage due to space constraints. From the results, we observe that PIA and DynamiCrafter, despite being trained on large-scale video datasets, generate unrealistic and disjointed motion that deviates considerably from the correct actions. This reveals the limitations of large-scale pretrained video generative models in animating human actions. DreamVideo, MotionDirector and LAMP finetune the models on a small set of videos containing the target actions. However, DreamVideo and MotionDirector exhibit obvious deviations from the reference images, indicating their difficulties in adapting motion to different reference images. LAMP shows smooth transition from the reference image but struggles with action fidelity, as seen in its rendering of the shoot dance, which exhibits disconnected or missing limbs, and its failure to generate the chopping wood action. In contrast, FLASH not only maintains smooth transition from the reference image but also realistically generates the intended actions that resemble real videos, demonstrating its effectiveness.
@@
Generalization to Diverse Images. To assess the generalization capability of FLASH beyond the HAA500 dataset, we test it on images sourced from the Internet and those generated by Stable Diffusion 3 [12]. As shown in Figure 5, FLASH successfully animates actors in unrealistic scenarios, such as (a) an astronaut running in place in a virtual space and (b) a cartoon character shooting a soccer ball. Additionally, FLASH can animate generated images, such as (c) a humanoid alien pouring water over his head and (d) two humanoid aliens hugging. The results highlight FLASH’s strong generalization ability across diverse reference images. The animated videos from different methods are on the Webpage.
Table 1: Quantitative comparison of different methods. The best and second-best results are bolded and underlined.
Method    Cosine RGB ($\uparrow$)    Cosine Flow ($\uparrow$)   CD-FVD ($\downarrow$) Text Alignment ($\uparrow$) Image Alignment ($\uparrow$) Temporal Consistency ($\uparrow$)
TI2V-Zero 68.91 50.07 1524.08 23.32 67.40 87.83
SparseCtrl 67.18 57.16 1584.19 21.74 61.01 88.28
PIA 69.91 60.04 1571.87 23.11 64.59 93.87
DynamiCrafter 78.82 64.09 1419.68 23.08 81.75 95.31
DreamVideo 67.71 63.40 886.70 23.93 65.93 93.50
MotionDirector 74.83 68.28 1099.32 21.91 74.82 95.53
LAMP 83.17 70.60 1240.89 23.17 78.57 93.81
FLASH 85.48 77.42 815.11 23.21 79.66 95.81
Automatic Evaluation Results. Following [91, 92, 29], we use three metrics based on CLIP [67]: Text Alignment, Image Alignment and Temporal Consistency, where higher scores indicate better performance. However, CLIP may have limited ability to capture fine-grained visual details [81], which may affect the accuracy of these metrics. To compare generated and real videos, we utilize Fréchet distance following [94]. Specifically, we adopt CD-FVD [15], which mitigates the content bias in the commonly used FVD [83], providing a better reflection of motion quality. A lower CD-FVD indicates better performance. However, since the distribution is estimated from a limited number of testing videos, CD-FVD may not fully capture the accurate distances. To provide more accurate similarity measurements between a generated video and its real counterpart with the same reference frame, we compute the cosine similarity between each generated video and its corresponding ground-truth video using the same reference image in HAA500. We calculate two metrics, Cosine RGB and Cosine Flow, using RGB frames and optical flows, respectively, where higher similarity values indicate better performance. For all metrics, we report the average results across all test videos. More details are described in Appendix A3.3.
Table 1 presents the quantitative comparison across six metrics, where FLASH achieves the best performance except in Text Alignment and Image Alignment. This indicates that FLASH excels in generating actions with the high temporal consistency and similarity to real action videos. For Text Alignment, TI2V-Zero and DreamVideo outperform FLASH, but both score significantly lower on Image Alignment, because they only generate text-aligned content but struggle to transition from reference images (see Figure 4). For Image Alignment, DynamiCrafter surpasses FLASH but performs considerably worse on other metrics, as it tends to replicate the reference images rather than generate realistic actions (see Figure 4).
Figure 5: Animated videos from FLASH using reference images from the Internet and generated by Stable Diffusion 3.
Performance on Common Actions. Though we mainly focus on uncommon actions, our experiments also include common actions. On drinking from a cup and hugging human, FLASH outperforms baselines by over 100 on CD-FVD and 7% on Cosine Flow, with comparable results on other metrics, suggesting that FLASH can also improve the animation quality of common actions through few-shot fintuning.
Table 2: Quantitative ablation studies on different components of FLASH. The best and second-best results are bolded and underlined.
Variant Strong Augmentation Motion Features Alignment Inter-frame Correspondence Alignment Detail Enhancement Decoder   Cosine RGB ($\uparrow$)   Cosine Flow ($\uparrow$)   CD-FVD ($\downarrow$) Text Alignment ($\uparrow$) Image Alignment ($\uparrow$) Temporal Consistency ($\uparrow$)
#1 83.80 68.06 1023.30 22.53 77.10 95.43
#2 83.98 70.61 932.92 22.48 76.72 94.91
#3 84.44 71.40 920.39 22.64 76.48 95.06
#4 84.32 71.72 938.21 22.70 76.31 94.84
#5 84.46 72.24 906.31 22.52 76.35 95.01
#6 84.51 72.33 908.39 22.77 76.22 95.31
@@ 4.2 Ablation Studies
Due to limited computational resources, we conducted ablation studies on four representative actions: sprint run, soccer shoot, canoeing sprint, and hugging human, which cover single-person actions, human-object interactions and human-human interactions and both small- and large-scale motions. We compare several model variants with incremental modifications, as outlined in Table 2. The quantitative and qualitative results are presented in Table 2 and Figure A5 in the Appendix, respectively. The animated videos are available on the Webpage.
@@
Comparing the quantitative results of Variant #2 with Variants #3, #4, and #5, we find that Variants #3, #4, and #5 improve CD-FVD, Cosine RGB, and Cosine Flow. Both Variants #3 and #4 enhance the Cosine RGB, and Cosine Flow. When combined, Variant #5 yields further enhancements in cosine similarity and a 25-point improvements in CD-FVD, with only a slight decrease in Image Alignment. Qualitative results also indicates improved fidelity in Variants #3, #4, and #5. For instance, motion in Variant #2 appears unrealistic in both actions. In the soccer shooting action, the person’s foot didn’t touch the soccer ball, and the leg appears disconnected in some frames. In the canoe paddling action, the hand positions on the paddle are inconsistent across frames. However, these issues are largely mitigated in Variants #3, #4, and #5. These results demonstrate the effectiveness of the Motion Alignment Module in learning accurate motion. By providing explicit guidance for learning appearance-general motion, the module directs the model toward generalizable motion, thereby improving the quality of the generated videos.
Comparing the quantitative results of Variant #5 and Variant #6, we observe that Variant #6 noticeably improves Text Alignment and Temporal Consistency and slightly improves Cosine RGB and Cosine Flow, without substantially affecting CD-FVD. Qualitatively, Variant #6 enhances some details (e.g., the soccer ball in certain frames in the soccer shooting action) and reduces noise in generated frames (see videos on the Webpage). These results suggest that the Detail Enhancement Decoder could compensate for detail loss or distortions in generated frames. Since the decoder operates on a frame-by-frame manner without considering inter-frame relations, it has minimal impact on motion patterns, leading to only slight effects on CD-FVD.
More Results. In Appendix A4.2, we demonstrate that the Motion Alignment Module improves motion quality across various few-shot settings (i.e., 8 or 4 videos per action class) and benefits from joint training across multiple action classes. Additionally, we conduct ablation studies on the hyperparameters of the Motion Alignment Module, the branches of the Detail Enhancement Module, and further evaluate the Detail Enhancement Module using DINO-V2 [62]. In Appendix A4.3, we show that FLASH surpasses the baselines on UCF Sports actions. Appendix A4.4 shows the applicability of FLASH to natural scene motion.
5 Conclusion
We tackle the challenge of few-shot human action animation and propose FLASH. We introduce the Motion Alignment Module to learn generalizable motion by forcing the model to reconstruct two videos with identical motion but different appearances using the same aligned motion patterns. Additionally, we employ the Detail Enhancement Decoder to enhance transition smoothness through multi-scale detail propagation. Experiments validate the effectiveness of FLASH in animating diverse images.